b) A solution of iron (III) chloride has a density of 1.118 g/mL. What volume of solution would have a mass of 1.75 kg?
c) How many moles of tungsten (Z=74) are there in 192.9g of tungsten? d) What is the mass of 6.77x10- nitrogen atoms?
Answer #1

b.) density of FeCl3=10118g/m1 Volume = ? , mas ss 1,75kg density = mass - volume volume = mass density = (1875*1000)g 10118gdolo mole X 60022x10 23 X2=66778129 Nitrage i mole = 6077X1024 6.022X1023x2 mass - 6 TIS 28 -6.022 mass = 607745x28 = 157.399Please give better rating